# Topological Phase Transition and Geometrical Frustration in Fourier Photonic Simulator

### v2

@@ Abstract
According to the convolution theorem [39], the integral of light intensity $I(\bm{u})$ multiplied by a modulation function $I_{M}(\bm{u})$ is equivalent to performing the convolution operation for all spins with the convolutional kernel, $\mathcal{F}^{-1}\left\{I_{M}(\bm{u})\right\}$ [30]. Since $I_{M}(\bm{u})$ modulates the light intensity in the Fourier plane, we call it the Fourier mask. When $\mathcal{F}^{-1}\left\{I_{M}(\bm{u})\right\}$ takes the form of summation over several delta functions, representing the interaction between a spin to others at discrete lattice sites, the corresponding Fourier mask can be described by
@@
where $\bm{r}^{(\alpha)}$ is the distance between coupling spin sites. The summation over $\alpha$ indicates that arbitrary range interactions can be achieved by a Fourier mask. Therefore, the Hamiltonian of the XY model can be directly obtained in experiment by an appropriate Fourier mask, which accelerates the sampling of spin configurations (see derivation in Supplementary Material [30]). In our experiment, the Fourier mask is linearly renormalized to values between 0 and 1, which allows experimental implementation with an intensity attenuator such as a metasurface [40].
In addition, this strategy is compatible with the free arrangement of spins on the SLM. It introduces a new programmable degree of freedom, enabling to customize special lattice geometries beyond square lattices. For instance, the honeycomb and kagome lattices can be constructed by introducing vacancies and displacements in arranging spins on the SLM. The vacancy means that the wavefront amplitude of incident light is set to zero at these sites, achieved by an additional intensity modulation or by directly implementing complex amplitude modulation. We then use the Markov chain Monte Carlo (MCMC) sampling algorithm, e.g., Metropolis-Hastings [41] and parallel tempering [42], to update the spin configurations and to study XY models across a range of lattice geometries. The sampling loop of FPS is shown in FIG. 1.
Ferromagnetic XY model on the square lattice. –We first construct the ferromagnetic XY model on the square lattice with NN interaction, as shown in FIG. 2(a). In this case, the $J_{\alpha}$ in Eq. 1 is $J\delta(\alpha-1)$ where $J$ is the interaction strength constant. Unlike the Ising model, the XY model has no phase transition from spontaneous symmetry breaking [43, 44], but possesses a BKT phase transition driven by interactions between topological defects [45]. These topological defects featured by vortices have significantly different dynamics above and below a phase transition temperature $T_{\text{BKT}}/J\approx 0.8922$ [46, 7]. Vortices with opposite charges, whose spins rotate clockwise and anti-clockwise, respectively, are tightly bound at low temperatures and dissociate above $T_{\text{BKT}}$ [47]. Conventionally, phase transition can be defined by appropriate order parameters characterized by symmetry breaking. However, a BKT phase transition has no symmetry breaking, such that $T_{\text{BKT}}$ cannot be related to order parameters.
The topological phase transition can be characterized from another perspective. The correlation function $G(\bm{r})=\langle\bm{S}(\bm{x}_{j})\cdot\bm{S}(\bm{x}_{j}+\bm{r})\rangle$ scales differently above and below the $T_{\text{BKT}}$ [49]. When $T>T_{\text{BKT}}$, $G(\bm{r})$ decays exponentially because the vortices are randomly distributed. When $T<T_{\text{BKT}}$, vortex pairs with opposite charges are tightly bound, such that the $G(\bm{r})$ decays in a power-law manner, indicating a quasi-long range order. In FIG. 2(c), we show $G(\bm{r})$ for the two different topological phases at $T/J=1.2$ and $0.8$. The exponential decay function fits better for the high-temperature phase, while the power-law decay function fits better for the low-temperature phase. Therefore, FPS can be used to precisely simulate XY models and reveal the complex topological phase transition.
Antiferromagnetic XY model on the triangular lattice.–Antiferromagnetic interactions can induce frustration in XY models on triangular lattices [50]. Spin pairs of nearest neighbors tend to align antiparallel due to the antiferromagnetic coupling. However, spins at three vertices within each triangular unit cannot simultaneously align antiparallel to each other. It is termed geometrical frustration and leads to novel phases and orders [31]. For instance, some researches claim that this geometrical frustration in antiferromagnetic triangular lattice induces an additional second-order phase transition, and the critical temperature is slightly higher than the BKT transition temperature [51, 50]. Nevertheless, many results support the existence of only one BKT phase transition [52, 53]. To experimentally validate this, we construct the antiferromagnetic XY model on a triangular lattice by FPS, and employ Metropolis-Hastings algorithm to investigate the frustration and phase transitions.
Figure 3: Simulating the antiferromagnetic XY model on a triangular lattice with $N=12\times 12$ spins. (a) The triangular lattice structure with the spin configuration of a ground state (upper penal), the interacting neighbors on the SLM (lower left), and the Fourier mask for NN interactions with periodic boundary conditions (lower right). (b) The specific heat ($C_{V}$) and helicity modulus ($Y$) as functions of temperature. $C_{V}(T)$ peaks at $T_{c}=0.48J$ indicating a second-order phase transition, while $Y(T)$ intersects $\tilde{Y}(T)=\frac{2}{\pi}T$ at $T_{\text{BKT}}\approx 0.43J$, denoting a BKT transition with $T_{\text{BKT}}<T_{c}$.
The triangular lattice is obtained on a square lattice by properly deforming and assigning interactions between each spin and its six neighbors. The corresponding deformed Fourier mask is shown in FIG. 3(a) (see the Supplementary Material [30] for details). We engineer the Fourier mask to include the NN interactions and the NNN interactions along the main diagonal of the square lattice. To distinguish the two different phase transitions, we use the specific heat $C_{V}=\frac{\beta^{2}}{N}\left(\langle\mathcal{H}^{2}\rangle-\langle\mathcal{H}\rangle^{2}\right)$ to characterize the second-order phase transition [50], while use the helicity modulus to characterize the BKT phase transition. As seen in FIG. 3(b), the specific heat has a peak at $T_{c}=0.48J$ and the intersection of $Y(T)$ and $\tilde{Y}(T)$ is at the $T_{\text{BKT}}\approx 0.43J$. Our results reveal two distinct phase transitions for the antiferromagnetic XY model on a triangular lattice, in which the second-order phase transition occurs at a slightly higher temperature than the BKT phase transition, consistent with high-accuracy numerical simulations [54]. Meanwhile, the geometrical frustration of the ground states is demonstrated by the angle between any two spin pairs, which is $32\pi/$ 3 within each triangular unit (see Fig. 3(a) for one of the ground states and Supplementary Material [30] for details). Therefore, FPS is a reliable tool for studying complex phase-transition behavior involving frustrations.
Antiferromagnetic $J_{1}$-$J_{2}$ XY model on the honeycomb lattice.–To demonstrate the ability of the FPS in lattice engineering, we further study the $J_{1}$-$J_{2}$ model on the honeycomb lattice. To construct a honeycomb lattice with the square lattice pixels of the SLM, we need to stretch the lattice, shift its sites and introduce vacancies, as shown in FIG. 4(a). The NN and NNN spin interactions $J_{1}$ and $J_{2}$ are realized by Fourier masks and the ratio $J_{2}/J_{1}$ is tunable from zero to infinity [32] by changing the superposition ratio of the two corresponding Fourier masks [16]. The competition between the NN and NNN interactions leads to geometrical frustrations and highly degenerate ground states at the zero temperature. These can be characterized in reciprocal space by the distribution of spiral wave vectors within the first Brillouin zone (1BZ) [32].
Figure 4: Searching the ground state of antiferromagnetic $J_{1}$-$J_{2}$ XY model on the honeycomb lattice with $N=96$ spins. (a) The lattice structure of the honeycomb lattice (upper left), the spin arrangement on the SLM (upper right), and the Fourier masks for NN interactions ($J_{1}$, lower left) and NNN interactions ($J_{2}$, lower right) with periodic boundary conditions. In building a honeycomb lattice from a square lattice, we set the light intensity in the dotted box areas in the SLM as zero. (b) The ground states in reciprocal space with different $J_{2}/J_{1}$ ratios. The black dashed lines mark the boundary of the 1BZ, while the white solid lines indicate the theoretically predicted positions of the peaks [32].
We experimentally found ground states with different $J_{2}/J_{1}$ ratios, as shown in FIG. 4(b). When $J_{2}=0$, only NN interaction is involved. In the ground state the nearest neighbor spin pairs are antiparallel. We obtain two sublattices with opposite spin orientations. Hence, the peak of the spiral wave vector is located at the center of the 1BZ ($\bm{\mathit{\Gamma}}$). It remains the case within the range $60\leq J_{2}/J_{1}\leq 1/$ 6 (FIG. 4(b1) and (b2)), indicating the Néel antiferromagnetic phase. When $21/6<J_{2}/J_{1}<1/$ 2, multiple peaks appear on a closed contour centered at the $\bm{\mathit{\Gamma}}$ point. The solid white lines in FIG. 4(b3) show the theoretical predictions of all the degenerate ground states. In the experiment the system will randomly relax to some of the ground states, corresponding to the separated peaks on these curves. Specifically, the peaks for $J_{2}/J_{1}=1/2$ are along white lines passing through the $\bm{M}$ points (FIG. 4(b4)). When we further increase the NNN interaction, $J_{2}/J_{1}>1/2$, the peaks appear on the closed contours around the $\bm{K}$ and $\bm{K}^{\prime}$ points (FIG. 4(b5)). When $J_{2}/J_{1}\rightarrow\infty$, the orientations of NN spin pairs in each sublattice differ by an angle $32\pi/$ 3, and the peaks in the 1BZ are shifted to the $\bm{K}$ and $\bm{K}^{\prime}$ points, which is a $32\pi/$ 3 ordered phase (FIG. 4(b6)). These experimental results are consistent with theoretical predictions [32], which demonstrates the generalizability of FPS for exploring statistical mechanics of spin systems with complex lattice geometries and spin interactions.
@@
Discussions.– Due to geometrical frustration, the energy landscape of the XY model is generally rough such that it takes a long time to obtain a sufficient number of samples for large-scale lattices. FPS directly obtains the Hamiltonian of spin configurations from the measured light intensity and updates it using MCMC algorithms, which significantly speeds up the sampling process. By incorporating the Fourier-mask technique into the FPS, we can employ a single Fourier mask to realize arbitrary range spin interactions. Consequently, the speed of calculation is nearly the same for any range of spin interactions, and the computational complexity is $\mathcal{O}(N)$, while most of the current local updating algorithms scale as $\mathcal{O}(N\mathrm{log}N)$ [33] (see the Supplementary Material [30] for detailed information). This feature enables the efficient calculation of large-scale XY models with complex long-range interactions [34], and can be further extended to simulate statistical models that can be mapped to XY models, including superfluids [46], superconductors [47], and arrays of Josephson junctions [55]. We also provide some results pertaining to long-range interactions featuring power-law decay on square lattices (see Supplementary Material [30]).
@@
The time consumption of FPS includes the response time of the SLM, the acquisition time of the camera, and the computation time of Hadamard multiplication and summation between $I$ and $I_{M}$. In our implementation of the FPS, the SLM is replaced by a digital micromirror device, which has a refresh rate of 20 kHz, significantly increasing the modulation speed of the FPS [56]. However, this step also introduces a large optical aberration caused by the industrial-grade chip of digital light processing, resulting in significant energy deviations. To address this issue, we develop an iterative retrieval approach to precompensate the overall aberration of FPS (see Supplementary Material [30]), enabling to precisely generate the Hamiltonian of the XY models. For further acceleration, a metasurface [40] and photodiode can be used to replace the camera to break the speed limit and reduce the overall time consumption. We can use the metasurface as the Fourier mask, such that the Hadamard multiplication and summation are automatically obtained from the light intensity given by the photodiode. We can further speed up the sampling procedure by using state-of-the-art machine learning algorithms to suggest optimized spin-flip policies [57], which can be used to solve optimization problems in real time [58].
In conclusion, we present a programmable Fourier photonic spin simulator to solve the XY models with various lattice geometries and complex spin interactions. Such FPS harnesses SLM to encode spins and the Fourier mask to obtain lattice geometries and spin interactions. This approach enables a lattice-size independent Hamiltonian calculation for the XY models. We experimentally demonstrated the FPS’s capabilities by solving problems involving topological phase transition and spin frustration in three distinct systems, i.e., the ferromagnetic square lattices, the antiferromagnetic triangular lattices, and the antiferromagnetic $J_{1}$-$J_{2}$ honeycomb lattices. In these models, we precisely determined the critical transition temperatures, directly observed the geometrically frustrated states, and identified the degenerate ground-state configurations. These findings validate the universality of the FPS for exploring various phases and critical behaviors across the XY models, and is promising to be extended to quantum spin systems [59, 60]. It also has the potential to solve continuous optimization problems, accelerate statistical learning algorithms, and establish large-scale optical neural networks.
This work was supported by the National Natural Science Foundation of China (Grant No. 12325412, 12088101, 12434020 and 11934011), National Key Research and Development Program of China (Grant No. 2024YFA1408900 and 2019YFA0308100), the Strategic Priority Research Program of Chinese Academy of Sciences (Grant No. XDB28000000), and the Fundamental Research Funds for the Central Universities.

### v3

@@ Appendix A End Matter
The Hamiltonian of the XY model is obtained by making integration of $I(\bm{u})I_{M}(\bm{u})$, where $I_{M}(\bm{u})$ is a Fourier mask [16] designed for specific spin interactions. According to the convolution theorem [59], the integral of light intensity $I(\bm{u})$ multiplied by a modulation function $I_{M}(\bm{u})$ is equivalent to performing the convolution for all spins with the kernel $\mathcal{F}^{-1}\left\{I_{M}(\bm{u})\right\}$ [34]. When $\mathcal{F}^{-1}\left\{I_{M}(\bm{u})\right\}$ takes the form of summation over several delta functions, representing the interaction between a spin to others at discrete lattice sites, the corresponding Fourier mask is
@@
where $\bm{r}^{(\alpha)}$ is the distance between interacting spin sites. $J_{\alpha}$ controls the ratio of various Fourier masks by weighted sum and the summation over $\alpha$ indicates that arbitrary range interactions can be achieved by a Fourier mask. Therefore, the Hamiltonian of the XY model can be directly obtained in experiment by a correspondingly designed Fourier mask, which accelerates the sampling of spin configurations (see derivation in [34]). In our experiment, the Fourier mask is linearly renormalized to values between 0 and 1, which allows experimental implementation with an intensity attenuator such as a metasurface [60].
@@
Computational complexity of the FPS–Due to geometrical frustration, the energy landscape of the XY model is generally rough such that it takes a long time to obtain a sufficient number of samples for large-scale lattices. FPS directly obtains the Hamiltonian of spin configurations from the measured light intensity and updates it using MCMC algorithms, which significantly speeds up the sampling process. By incorporating the Fourier-mask technique into the FPS, we can employ a single Fourier mask to realize arbitrary range spin interactions. Consequently, the speed of calculating the Hamiltonian is nearly the same for any range of spin interactions, and the computational complexity is $\mathcal{O}(N)$, while most of the current local updating algorithms scale as $\mathcal{O}(N\mathrm{log}N)$ [37] (see [34] for detailed information). This feature enables efficient calculation of large-scale XY models with complex long-range interactions [38], and can be further extended to simulate statistical models that can be mapped to XY models, including superfluids [49], superconductors [43], and Josephson junctions arrays [61]. We also provide some results on long-range interactions featuring power-law decay in square lattices (see [34]).
@@
The time consumption of FPS includes the response time of the SLM, the acquisition time of the camera, and the computation time of Hadamard multiplication and summation between $I$ and $I_{M}$. In our implementation of the FPS, the SLM is replaced by a digital micromirror device, which has a refresh rate of 20 kHz, significantly increasing the modulation speed of the FPS [62]. However, this step also introduces a large optical aberration caused by the industrial-grade chip of digital light processing, resulting in significant energy deviations. To address this issue, we develop an iterative retrieval approach to precompensate the overall aberration of FPS (see [34] for the algorithm details and the results of the aberration compensation), enabling to precisely obtain the Hamiltonian of the XY models. For further acceleration, a metasurface [60] and photodiode can be used to replace the camera to reduce the overall time consumption. We can further speed up the sampling procedure by using state-of-the-art machine learning algorithms to suggest optimized spin-flip policies [63], which can be used to solve optimization problems in real time [64].